# Position: On-Premises LLM Deployment Demands a Middle Path: Preserving Privacy Without Sacrificing Model Confidentiality

### v2

@@ B.9 Effectiveness of distillation difficulty
DD is effective in larger models, with a clear negative correlation between DD and average distillation ratios. As shown in Figure 9 (a), the Pearson coefficient for Llama2-7B consistently remains below -0.80, reaching as low as -0.98. Similar trends are observed in other models with varying architectures and sizes, confirming DD as a reliable predictor of distilled model performance and demonstrating the effectiveness of SOLID. Furthermore, the complete Pearson and Spearman results, presented in Table 11, reveal a negative correlation between RS and the average distillation ratio. For instance, in Llama2-7B, both the Pearson and Spearman coefficients fall below -0.77. These trends hold across models of different architectures and sizes. Additionally, Figure 10 displays scatter plots illustrating the relationship between $\Delta$ADR and Distillation Difficulty($\uparrow$)s for four models, with corresponding Pearson and Spearman correlation coefficients. The Distillation Difficulty($\uparrow$)s, obtained from Section 5.3, show a clear inverse relationship: an increase in $\Delta$ADR leads to a decrease in model scores across all models analyzed. This trend is consistently supported by strong negative correlations, with the most significant drop in model scores seen in Phi2-2.7B, where $\Delta$ADR shows the most substantial negative correlation.
DD is ineffective in smaller OPT model, with notable inconsistencies with ADRs. As shown in Figure 9 (b), the negative correlation between DD and ADR in OPT-350M is weak (coefficients above -0.33), showing its unsuitability for predicting distillation performance. Additionally, optimal security is achieved by protecting the middle layers rather than the initial or output layers, making SOLID unable to identify the smallest secured set.
@@
Model Rsn. Read. Knl. Code & Math Gen. Avg.
Llama2-7B -0.83 |||| -0.97 -0.77 |||| -0.96 -0.83 |||| -0.95 -0.85 |||| -0.90 -0.82 |||| -0.93 -0.80 |||| -0.98
Mistral-7B -0.83 |||| -0.89 -0.82 |||| -0.91 -0.82 |||| -0.94 -0.78 |||| -0.95 -0.76 |||| -0.87 -0.87 |||| -0.92
Phi-2 -0.93 |||| -0.96 -0.84 |||| -0.96 -0.84 |||| -0.87 -0.84 |||| -0.80 -0.84 |||| -0.84 -0.87 |||| -0.95
Phi-1.5 -0.86 |||| -0.97 -0.78 |||| -0.94 -0.83 |||| -0.94 -0.90 |||| -0.80 -0.84 |||| -0.89 -0.80 |||| -0.94
Figure 10: Correlation Analysis of $\Delta$ADR and Distillation Difficulty Across Different Models.
@@ C.1 Comparison in two semi-open Llama2-70B
Table 13: Distillation Performance of Llama2-70B under Different Secured Layers
@@ C.2 Evaluation Results under FT-all attack
Table 14: Evaluation results of Llama2-70B under FT-all attack
@@
Table 15: Evaluation results of 7B models under FT-all attack
@@
Table 16: Evaluation results of small models under FT-all attack
@@ C.5 Comparison in Distillation Attack Strategies
Table 23: Distillation Performance of SOLID under FT-Secure attacks.
@@
Table 24: Distillation Performance of SOLID under SEM attacks.
@@ C.6 Comparison in Distillation datasets scales
Table 27: Evaluation Results of Llama2-7B under Different Secured Layers (Part2). * indicates the fully secured model.
@@
Table 33: Customization Performance under Different Secure Sets
@@ C.9 Limitation on OPT-350M
[…]

### v3

@@ B.9 Effectiveness of distillation difficulty
[…]
Model Rsn. Read. Knl. Code & Math Gen. Avg.
Llama2-7B -0.83 $|$ -0.97 -0.77 $|$ -0.96 -0.83 $|$ -0.95 -0.85 $|$ -0.90 -0.82 $|$ -0.93 -0.80 $|$ -0.98
Mistral-7B -0.83 $|$ -0.89 -0.82 $|$ -0.91 -0.82 $|$ -0.94 -0.78 $|$ -0.95 -0.76 $|$ -0.87 -0.87 $|$ -0.92
Phi-2 -0.93 $|$ -0.96 -0.84 $|$ -0.96 -0.84 $|$ -0.87 -0.84 $|$ -0.80 -0.84 $|$ -0.84 -0.87 $|$ -0.95
Phi-1.5 -0.86 $|$ -0.97 -0.78 $|$ -0.94 -0.83 $|$ -0.94 -0.90 $|$ -0.80 -0.84 $|$ -0.89 -0.80 $|$ -0.94
@@
[…]
B.10 Adversarial Attack
In this section, we provide a detailed comparison of SOLID and SAP-DP in their effectiveness against three types of black-box adversarial attacks on the Llama2-7B model. The attacks considered include Membership Inference Attacks (MIA), Attribute Inference Attacks (AIA), and Prompt Injection Attacks (PIA).
Membership Inference Attack (MIA): This attack aims to determine whether a specific data point was included in the training dataset of the model. Attackers utilize model outputs to infer membership status, potentially exposing sensitive information about the training data (Fu et al., 2023; Chen and Pattabiraman, 2024). We conducted our experiment following SPV_MIA  [^9] https://github.com/wjfu99/MIA-LLMs, which provides a robust framework for assessing model vulnerabilities. We focus on the AUC scores for SPV-MIA against semi-open models across Ag News datasets (Zhang et al., 2016).
Attribute Inference Attack (AIA): In this scenario, the adversary attempts to infer specific attributes of training data based on the model’s outputs. This can lead to privacy breaches, particularly when sensitive attributes are involved (Staab et al., 2023; Li et al., 2024b). We conducted our experiments following the methodology outlined in Staab et al. (2023)  [^10] https://github.com/eth-sri/llmprivacy and evaluated the top-3 accuracy on the PersonalReddit (PR) Dataset.
Prompt Injection Attack (PIA): This attack manipulates input prompts to coerce the model into producing desired outputs that may compromise the integrity or security of the system (Zhao et al., 2024; Xu et al., 2024b). In our experiment, we follow AutoDAN  [^11] https://github.com/SheltonLiu-N/AutoDAN, which can automatically generate stealthy jailbreak prompts by the carefully designed hierarchical genetic algorithm.We evaluate the effectiveness of these prompts using the keyword-based attack success rate (ASR), which measures the presence of predefined keywords in responses generated LLMs. For gold standard, LED  [^12] https://github.com/ledllm/ledllm, significantly enhances the security of LLMs against prompt injection attacks (PIA), reducing the ASR to 0.
Limited Defense against Adversarial Attack. We compare SOLID and SAP-DP in defending against three black-box adversarial attacks on Llama2-7B. Specifically, we apply the membership inference (Fu et al., 2023) (MIA), attribute inference (Staab et al., 2023) (AIA), and prompt injection (Liu et al., 2023) (PIA) attacks to the semi-open models produced by SAP-DP and SOLID. As shown in Table 14, we observe that SAP-DP outperforms SOLID across all three attacks, but still performs worse than the gold standard. This is because SOLID does not introduce additional output perturbation and thus provide limited defense against black-box adversarial attacks. Details can be found in Appendix B.10.
[…]
Gold Std. 58.0 43.9 0.00
SCARA 72.3 85.0 26.5
SAP-DP 72.2 83.9 24.9
Table 14: Performance of SCARA defending adversarial attacks. $\downarrow$ indicates the smaller the better.
@@ C.2 Evaluation Results under FT-all attack
Table 17: Evaluation results of Llama2-70B under FT-all attack
@@
Table 18: Evaluation results of 7B models under FT-all attack
@@
Table 19: Evaluation results of small models under FT-all attack
@@ C.5 Comparison in Distillation Attack Strategies
Table 26: Distillation Performance of SOLID under FT-Secure attacks.
@@
Table 27: Distillation Performance of SOLID under SEM attacks.
@@ C.6 Comparison in Distillation datasets scales
Table 30: Evaluation Results of Llama2-7B under Different Secured Layers (Part2). “*” indicates the fully secured model.
@@
Table 36: Customization Performance under Different Secure Sets
@@ C.9 Limitation on OPT-350M
To investigate the limitations of SOLID, we calculate the Distillation ratio of each secure-source set within the smaller model, OPT-350M (Zhang et al., 2022) with only 350M parameters. We set the secure-source set size to 2 and subsequently calculate $\Delta\text{ADR}$s for each secure-source set. The detailed results are shown in Figure 45.